40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5511     Entry time: Thu Sep 22 01:05:28 2011
Author: Katrin 
Type: Update 
Category: Green Locking 
Subject: New modulation frequency (Y arm) 

[Kiwamu / Katrin]

 

On Wednesday, the green light was locked to the Y arm cavity.

Modulation frequency was changed from 279kHz to 178875Hz. The amplitude was changed from 10Vpp to 0.01Vpp to achieve a modulation index of 0.38. The modulation frequency was changed to minimize AM. With the new modulation frequency the laser light could still locked to the cavity.

The signal of the LO and the photodiode are multiplied by a ZAD-8 mini circuit mixer (Level 7). This mixer requires LO input is +7dBm = 1.4Vpp. Thus we put a 36dB attenuator between the LO and the PZT at the laser. PDH error signal shows lots of peaks that are most likely higher order sidebands. Thus, the next step is to work on the low-pass filter. However the SNR of the error signal has improved with the new modulation frequency. With the old mod. frequency the PDH signal was 4mVpp and the noise floor was 2mVpp.

Phase between the photodiode signal and LO is shifted by about 10 degrees. Step two is to work on a phase shifter.

 

 

ELOG V3.1.3-